4 t2= The 2 is squared... so 4t 2=16?
Found 2 solutions by stanbon, sinadone:
Answer by stanbon(75887)

4t^2=16
Divide thru by 4 to get
t^2=4
Take the square root of boh sides to get:
t=2 or t=-2
Cheers,
Stan H.

Answer by sinadone(8)

4t^2=16
Divide by 4:
t^2=4
Then take the square root of t^2 and of 4.
which leaves you with: t=2
Whenever you want to recheck your answer, just plug it back in.
4(2)^2 (should equal) = 16
Then you know you're right!